OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by the given curves about the specified line. sketch the region and the solid. y=1-x^2, y=0; about the x-axis

OpenStudy (anonymous):

about x axis

OpenStudy (anonymous):

\[ V = \pi \int\limits_{a}^{b} y^2 dx \]

OpenStudy (anonymous):

\[ V= \pi \int\limits_{a}^{b} ( 1-2x^2 +x^4 ) dx \]

OpenStudy (anonymous):

where a<b , and a and b are boundaries on x

OpenStudy (anonymous):

y=0 is the x axis :\

OpenStudy (anonymous):

the region is not defined

OpenStudy (anonymous):

if x=0 would answer be c + (8pi)/15

OpenStudy (anonymous):

:|

OpenStudy (anonymous):

what u said makes absolutely no sense, and you included a constant in a definate integral

OpenStudy (anonymous):

what would the answer be if boundaries were x=0 and y= 0, and Y= 1- X^2 revolving around x axis

OpenStudy (anonymous):

8pi/15 I guess, without the c :|

OpenStudy (anonymous):

u guess? u didnt try to work it out did u. just seein ghow rusty i am

OpenStudy (anonymous):

so what is the answer on my question